OpenStudy (anonymous):

x = t^2 + 3 y = 2t Eliminate the parameter and write a rectangular equation

OpenStudy (anonymous):

substitute

OpenStudy (anonymous):

substitute what?

OpenStudy (anonymous):

t = y/2

OpenStudy (anonymous):

(y/2)^2 + 3 ?

OpenStudy (anonymous):

x = that

OpenStudy (anonymous):

thats what the answer is ?

OpenStudy (anonymous):

you tell me... do you know the definition of rectangular equation? If not, look it up to be sure

OpenStudy (anonymous):

so yes , thats the answer..

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

(y/2)^2 + 3 is that the answer, or no ?

OpenStudy (anonymous):

i'm too lazy to say... you're too lazy to look

OpenStudy (anonymous):

but i know the answer and you need it

OpenStudy (anonymous):

and all you have to do is look at what i sent

OpenStudy (anonymous):

so you don't need me any more just get off your butt and do something

OpenStudy (anonymous):

if you don't want to help just don't reply. its that easy. dont need to be mean about it.

OpenStudy (anonymous):

i'm not, but you're not asking for help... just answers

OpenStudy (anonymous):

di you look at what i sent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what part is confounding you?

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

what is a parametric equation?

OpenStudy (anonymous):

just help me understand , i replaced the t with 2/y

OpenStudy (anonymous):

basically the question i gave you

OpenStudy (anonymous):

did you see them do the same in the attachement? did you look at page 2, ex. 2?

OpenStudy (anonymous):

so the answer is 3+y^2/2

OpenStudy (anonymous):

no... t = y/2. so t^2 = (y/2)^2 not t = y^2/2 they're different

OpenStudy (anonymous):

so its 3 + (y/2) ^2

OpenStudy (anonymous):

or 3 + y^2 / 4